OpenStudy (anonymous):

Solve the equation: x2 - 12x = -36

OpenStudy (theviper):

Use Quadratic Formula

OpenStudy (anonymous):

Just pull over the 36 to the other side: So you get x^2 -12x+36 = 0 Then use quadratic formula.

OpenStudy (ryan123345):

try and make it look like Ax^2+Bx+C=0

OpenStudy (anonymous):

ok i understand x^2 -12x+36 = 0 but i have to use factoring to solve it bcs i have never learned quads yet

OpenStudy (anonymous):

x2-12x= -36 x2-12x+36=0 (x-6) (x-6)=0

OpenStudy (anonymous):

now solve for x which is 6

OpenStudy (anonymous):

use FOIL to go back and look if the factoring was correct

OpenStudy (anonymous):

thank you